OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 5(x + 6)2(x - 6)3

OpenStudy (anonymous):

Is that\[f(x) = 5(x+6)^2(x-6)^3\]?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

set y = 0, thus \(\bf f(x) = 5(x + 6)^2(x - 6)^3\implies 0=5(x + 6)^2(x - 6)^3\) so... what are the zeros?

OpenStudy (anonymous):

-6 and 6?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

soo -6 would have multiplicity of 3? and 6 of 2? wait is 4 a zero too?

OpenStudy (jdoe0001):

the multiplicity it's just \(\bf 0=5(x + 6)^\color{blue}{2}(x - 6)^\color{red}{3}\implies 0=5\color{blue}{(x + 6)(x + 6)}\color{red}{(x - 6)(x - 6)(x - 6)}\)

OpenStudy (anonymous):

okay so no 4

OpenStudy (jdoe0001):

yes, the multiplicity, or the times they REPEAT is 2 and 3 respectively

OpenStudy (anonymous):

thankyou :)

OpenStudy (jdoe0001):

yw